1.    Consider the random quantity X, with PDF

   fx(x) = {4/3(2x-x^2), 0<x<2
            0,         otherwise. 

 

i.                  Check that this is a valid PDF (it integrates to 1).

ii.                Calculate the expectation and variance of X.

iii.              Evaluate the cumulative distribution function.

iv.               What is the median of this distribution? 



1
Expert's answer
2022-02-21T15:50:15-0500

Given that,

"f(x)={3\over4}(2x-x^2), 0\lt x\lt 2\\0,elsewhere"


"a)" verify that it is a pdf.

"\displaystyle\int^2_0f(x)dx=\displaystyle\int^2_0{3\over4}(2x-x^2)dx ={3\over4}(x^2-{x^3\over3})|^2_0={3\over4}(4-{8\over3})={3\over4}\times{4\over3}=1.0"

Therefore, it is a valid pdf.


"b)"

"i)" Expectation of "X"

It is given as,

"E(x)=\displaystyle\int^2_0 xf(x)dx=\displaystyle\int^2_0{3\over4}(2x^2-x^3)dx={3\over4}({2\over3}x^3-{x^4\over4})|^2_0=1"

"ii)" Variance of "X"

The variance of "X" is given as,

"Var(x)=E(x^2)-(E(x))^2"

We determine,

"E(X^2)=\displaystyle\int^2_0 x^2f(x)dx=\displaystyle\int^2_0{3\over4}(2x^3-x^4)dx={3\over4}({x^4\over2}-{x^5\over5})|^2_0={6\over5}"

Therefore,

"var(x)={6\over5}-1^2={1\over5}"


"c)"

The cumulative distribution for "X" is,

"F(x)=\displaystyle\int^x_0f(u)du=\displaystyle\int^x_0{3\over4}(2u-u^2)du={3\over4}(u^2-{u^3\over3})|^x_0={3\over4}(x^2-{x^3\over3})"

It can be written as,

"F(x)={3\over4}(x^2-{x^3\over3}), 0\lt x\lt 2\\0,elsewhere"


"d)"

To find the median, we find a value "y" such that,

"\displaystyle\int^y_0f(x)dx=0.5"

So,

"\displaystyle\int^y_0{3\over4}(2x-x^2)dx={3\over4}(x^2-{x^3\over3})|^y_0={3\over4}(y^2-{y^3\over3})=0.5"

We need to solve for the value of "y"

"{3\over4}(y^2-{y^3\over3})=0.5\implies (y^2-{y^3\over3})={2\over3}\implies y^2(1-{y\over3})={2\over3}\implies y=0.82 \space or\space 1"

The value of the median must be around the middle of the range (0,2). So the most appropriate value of the median is y=1